### v1

# Efficient Neural Video Representation via Structure-Preseving Patch Decoding

Abstract
Implicit Neural Representations (INRs) have attracted significant interest for their ability to model complex signals by mapping spatial and temporal coordinates to signal values. In the context of neural video representation, several decoding strategies have been explored to balance compactness and reconstruction quality, including pixel-wise, frame-wise, and patch-wise methods. Patch-wise decoding aims to combine the flexibility of pixel-based models with the efficiency of frame-based approaches. However, conventional uniform patch division often leads to discontinuities at patch boundaries, as independently reconstructed regions may fail to form a coherent global structure. To address this limitation, we propose a neural video representation method based on Structure-Preserving Patches (SPPs). Our approach rearranges each frame into a set of spatially structured patch frames using a PixelUnshuffle-like operation. This rearrangement maintains the spatial coherence of the original frame while enabling patch-level decoding. The network learns to predict these rearranged patch frames, which supports a global-to-local fitting strategy and mitigates degradation caused by upsampling. Experiments on standard video datasets show that the proposed method improves reconstruction quality and compression performance compared to existing INR-based video representation methods.
Index Terms:
Implicit neural representation, video compression, video representation.
I Introduction
Implicit Neural Representations (INRs) have emerged as a compact, resolution-independent alternative to conventional explicit representations for modeling complex signals. By expressing data as continuous functions of spatial and temporal coordinates through neural networks, INRs significantly reduce memory usage. They also support desirable properties such as infinite-resolution rendering and smooth interpolation. These advantages make them well-suited for modeling fine-grained signals in domains such as novel view synthesis [1, 2, 3, 4], image representation [5, 6, 7, 8], and increasingly, video representation [9, 10, 11].
Figure 1: Comparison of decoding strategies for neural video representation. (a) Coordinate-based: predicts pixels from spatio-temporal coordinates. (b) Frame-based: generates entire frames from embeddings. (c) Patch-based: reconstructs frames from uniformly divided patches. (d) SPP-based (ours): rearranges pixels into spatially aligned patches for structure-preserving, global-to-local reconstruction.
INR-based methods for neural video representation can be categorized by their decoding granularity: coordinate-based (pixel-wise) [12, 13], frame-based (image-wise) [9, 10, 11], and patch-based approaches [14, 15]. Each approach presents a different trade-off between flexibility, computational cost, and structural consistency. Coordinate-based approaches predict individual pixel values based on spatial and temporal inputs, offering unmatched resolution flexibility and the ability to capture intricate local variations. However, they are often computationally intensive and limited in capturing global context due to their inherently localized nature. Frame-based approaches, such as Neural Representations for Videos (NeRV) [9], mitigate this issue by predicting entire frames from compact embeddings, such as time indices or learned content vectors. These models benefit from faster inference and improved global structure modeling, but tend to struggle with high-frequency detail due to upsampling artifacts and spectral bias (favoring low-frequency components) [16, 17, 18, 19]. Patch-based methods lie between these extremes. By reconstructing each frame from smaller subregions, they aim to combine the detail-preserving nature of pixel-wise decoding with the structural coherence of frame-wise decoding. Nonetheless, conventional patch partitioning via uniform spatial division can disrupt continuity across patch boundaries. Independently reconstructed patches often fail to align perfectly, resulting in visible seams and reduced visual quality. This limitation motivates the exploration of patch-based methods that maintain global spatial coherence without sacrificing local detail fidelity.
Figure 2: Toy experiment comparing the fitting performance of one-dimensional signals. Top row (left to right): ground-truth signal, point-wise fitting, and whole-wise fitting. Bottom row (left to right): training loss curves, patch-wise fitting, and the proposed Structure-Preserving Patch (SPP)-wise fitting. While conventional patch-wise approaches suffer from boundary discontinuities, the proposed SPP-wise approach maintains structural continuity and enables accurate local fitting.
To overcome these limitations, we propose the Structure-Preserving Patch (SPP) method for neural video representation. Instead of independently decoding spatially divided patches, our method rearranges each frame into multiple sub-images through a pixel redistribution operation inspired by PixelUnshuffle. This operation preserves the relative spatial structure of the original frame across all patches. The model is trained to predict these structurally aligned patch images, enabling it to first capture global structure and then refine local content. This global-to-local fitting process reduces the risk of discontinuities at patch boundaries and leads to improved visual quality. We validate the effectiveness of our method through experiments on several benchmark video datasets. The results show that our structure-aware patch decoding improves reconstruction quality and compression performance compared to prior INR-based video representation models.
II Related works
II-A Implicit Neural Representation
The central concept of INR is to model a signal as a continuous function implemented by a neural network. Formally, a signal is defined as:
$y=\Phi_{\theta}(x),$ (1)
where $x$ is a continuous coordinate, $y$ is the signal value at that location, and $\theta$ denotes the learnable parameters of the network. The function $\Phi_{\theta}$ is trained to fit a given target signal through optimization. A seminal work in this field is Neural Radiance Fields (NeRF) [1], which models a 3D scene as a radiance field defined over spatial coordinates and viewing directions. NeRF demonstrated that photo-realistic view synthesis is possible by optimizing a fully connected network to predict color and density values along camera rays. Although computationally demanding, NeRF highlights the potential of neural functions to model complex continuous structures without relying on explicit mesh or voxel representations. To enhance the capability of INRs in modeling high-frequency details, SIREN [12] introduced a sinusoidal-activated network capable of modeling high-frequency content with improved gradient flow. By leveraging periodic activation functions, SIREN reduces the spectral bias inherent in ReLU-based networks and improves the network’s ability to capture fine details and textures. These foundational works have inspired the application of INRs to spatio-temporal domains, including video data. Their success shows that compact and continuous neural representations can be extended from static to dynamic signals, motivating recent research into INR-based video modeling.
Figure 3: Overview of the proposed Structure-Preserving Patchs (SPPs) based video representation framework. (a) Overall architecture. An input frame $I_{t}$ is encoded into a frame-level embedding $z_{t}$, which is then decoded into patch images $I_{t,i}$ using the temporal index $t$ and patch index $i$. The full frame is reconstructed by spatially rearranging the decoded patches. (b) Decoder details. Temporal and patch indices are embedded using positional encodings and MLPs. Early decoder layers are conditioned on $t$ to capture global structure, while later layers use $i$ to refine local patch details. (c) Patch construction via SPPs. Each frame is rearranged into multiple patch images (e.g., $I_{t,0}$, $I_{t,1}$, $I_{t,2}$, $I_{t,3}$) that preserve spatial consistency. This design supports a smooth global-to-local reconstruction process.
II-B Neural Video Representation
INRs have recently been explored as an alternative paradigm for video representation and compression. A representative example is NeRV [9], which encodes an entire video in the weights of a neural network that maps a frame index to its corresponding RGB frame. Unlike conventional INR with coordinate mapping, NeRV treats the video as a global function of learned time, effectively embedding the video content into the network parameters. Video decoding then requires only a forward pass for a given frame index. This formulation means that compressing the network (through techniques like weight pruning or quantization) effectively compresses the video. Building on this idea, E-NeRV [10] improves NeRV by disentangling spatial and temporal components. It learns separate embeddings for each domain and combines them during decoding via Adaptive Instance Normalization (AdaIN) [20], thereby improving both representational efficiency and compression performance. HNeRV [11] further builds on this concept by replacing content-agnostic embeddings derived from the frame index with content-adaptive embeddings. This adjustment allocates greater modeling capacity to complex or high-frequency frames. More recently, BoostingNeRV [21] introduced a conditional decoding framework that can enhance various existing INR-based video models. It modulates intermediate features using a temporal affine transformation based on frame indices, enabling improved reconstruction performance across different architectures.
In addition, patch-wise and hybrid methods have been proposed to improve reconstruction quality and scalability. For instance, PS-NeRV [14] reconstructs videos on a patch-by-patch basis instead of generating full frames in one shot. This patch-wise process offers flexibility in handling higher resolutions and can distribute computational load, albeit with the risk of patch boundary artifacts. To better capture video dynamics, DNeRV [22] adopts a dual-stream architecture. One stream models the static background, while the other captures inter-frame differences, allowing motion to be represented implicitly. Several other approaches introduce hierarchical model structures [23, 24, 25], incorporating explicit residual connections [26, 27], or emphasizing the preservation of frequency-domain characteristics [28, 29, 30]. Each of these innovations addresses specific shortcomings of basic INR video models, such as temporal flicker and poor high-frequency reconstruction. Our work aligns with this line of research, focusing on patch-based representation. We specifically aim to preserve global image structure in a patch-wise model to avoid the common pitfalls of patch division.
III Proposed method
III-A Overview
We propose a neural video representation method based on Structure-Preserving Patches (SPPs), as illustrated in Fig. 3. In this framework, each frame is decoded at the patch level. Unlike conventional patch-based methods, however, the patches are constructed to retain the original spatial layout of the frame. The key idea is to avoid treating patches as independent tiles and to generate patch images in a way that preserves the layout of the frame. These patch images tend to share similar structures and contain redundant information, which the network can exploit to improve representation efficiency. The decoding process is designed to follow a global-to-local fitting strategy. The model first learns a coarse global representation of the frame and then progressively refines local details within each patch. This hierarchical fitting process improves the network’s ability to reconstruct both global context and fine details. By preserving spatial continuity across patch boundaries, our method mitigates the quality degradation commonly observed in naive patch-based upsampling or reconstruction.
III-B Motivation
To highlight the importance of the global-to-local strategy, we conduct a simple one-dimensional signal-fitting experiment (Fig. 2). We compare four strategies: (1) Point-wise, where the signal is fitted at each coordinate independently; (2) Patch-wise, where the signal is divided into fixed-length segments and each segment is fitted separately; (3) Whole-wise, where the entire signal is fitted as a single input; and (4) SPP-wise, our proposed approach, which fits a rearranged version of the signal based on fixed-position samples, mimicking the structure-preserving patch mechanism. The results show that the whole-wise and SPP-wise strategies achieve superior fitting accuracy by preserving the overall structure of the signal. In contrast, point-wise and patch-wise methods prioritize local fitting and fail to maintain global consistency. These observations suggest that modeling the global structure first, followed by local refinement, is essential for accurate signal reconstruction. This insight forms the foundation of our global-to-local decoding framework used for video representation.
Figure 4: Visualization of reconstructed frames from different sequences. From top to bottom: “drift-straight” and “mallard-fly” from the DAVIS dataset, and sequences #10 and #12 from the MCL-JCV dataset. Our structure-preserving patch (SPP) based method maintains sharp edges and consistent spatial structure across frames, demonstrating effective reconstruction quality across diverse scenes.
III-C Structure-Preseving Patch Decoding
Our framework predicts each video frame at the patch level using a neural network that follows a global-to-local decoding strategy, as illustrated in Fig.3. For each input frame $I_{t}$ at time $t$, we divide it conceptually into $N$ spatially structured patches. To implement this, we apply a pixel rearrangement operation, similar to PixelUnshuffle, which transforms the frame into $N$ patch images: $I_{t,0},I_{t,1},\dots,I_{t,N-1}$. This rearrangement preserves the original spatial relationships across patches. The encoder, implemented using ConvNeXt blocks [31, 11], first processes the full frame $I_{t}$ to produce a compact frame-level embedding $z_{t}$. This embedding is passed to a decoder, which generates each patch image $I_{t,i}$ under the influence of two indices: the temporal index $t$ and the patch index $i$. In the decoder, early layers are modulated only by $t$, allowing the model to capture the global structure shared across all patches in the frame. Later layers are additionally conditioned on $i$, enabling the network to refine local details specific to each patch. Once all patch outputs $I_{t,i}$ are obtained, we reconstruct the full-resolution frame $\hat{I}_{t}$ by spatially rearranging them. This hierarchical conditioning mechanism allows the decoder to first build a coherent global structure, then specialize to local content, resulting in improved reconstruction quality. In addition, the structure-preserving nature of our patch generation helps avoid boundary discontinuities typically associated with naive patch-based decoding. The decoder is composed of multiple blocks $B_{n}$, each of which integrates the SNeRV block and the TAT Residual Block introduced in BoostingNeRV [21].
TABLE I: Comparison of reconstructed video quality
[…]
Method PSNR $\uparrow$ MS-SSIM $\uparrow$ LPIPS $\downarrow$
NeRV 28.60 0.8811 0.4150
HNeRV 30.69 0.9146 0.3476
Boost 33.53 0.9604 0.2674
Ours 34.23 0.9643 0.2539
III-D Loss Function
For each patch $i$ of a frame, we define a composite loss $\mathcal{L}_{i}$ that accounts for spatial accuracy, perceptual quality, and frequency-domain consistency:
$\mathcal{L}_{i}=w_{i}(\alpha\mathcal{L}_{1}(x_{i},\hat{x}_{i})+\beta\mathcal{L}_{\text{MS-SSIM}}(x_{i},\hat{x}_{i}))+\mathcal{L}_{\text{freq}}(x_{i},\hat{x}_{i}),$ (2)
$\mathcal{L}_{\text{freq}}(x,\hat{x})=\mathcal{L}_{1}(\text{FFT}(x),\text{FFT}(\hat{x})),$ (3)
[…]
$w_{i}=\frac{\sum_{j\neq i}\|x_{i}-x_{j}\|_{1}}{\sum_{k}\sum_{j\neq k}\|x_{k}-x_{j}\|_{1}+\epsilon},$ (4)
where $\epsilon$ is a small constant to avoid division by zero. This formulation normalizes the weights across all patches. The model focuses more on learning consistent structures while maintaining flexibility for difficult regions.
TABLE II: Comparison of reconstructed video quality
[…]
Method PSNR $\uparrow$ MS-SSIM $\uparrow$ LPIPS $\downarrow$
NeRV 31.64 0.9217 0.4126
HNeRV 33.47 0.9417 0.3571
Boost 35.60 0.9638 0.3039
Ours 35.94 0.9654 0.2936
[…]
Method Beauty Bosphorus HoneyBee Jockey ReadySetGo ShakeNDry YachtRide Average
NeRV 32.93 / 0.8843 33.09 / 0.9288 37.08 / 0.9780 31.06 / 0.8767 24.66 / 0.8205 32.67 / 0.9264 27.75 / 0.8573 31.22 / 0.8937
HNeRV 33.18 / 0.8876 17.57 / 0.5885 38.97 / 0.9838 31.75 / 0.8884 25.09 / 0.8358 33.73 / 0.9337 27.95 / 0.8546 29.44 / 0.8470
Boost 33.70 / 0.8980 35.81 / 0.9631 39.52 / 0.9852 33.84 / 0.9253 27.72 / 0.9070 35.55 / 0.9532 29.03 / 0.8967 33.45 / 0.9311
Ours 33.82 / 0.8992 36.08 / 0.9644 39.60 / 0.9854 34.21 / 0.9249 28.03 / 0.9054 35.96 / 0.9584 29.33 / 0.8945 33.70 / 0.9312
IV Experiment
TABLE IV: PSNR over training epochs on the Bunny video
[…]
Epoch 50 100 150 200 250 300
NeRV 26.95 29.40 30.43 30.96 31.14 31.33
HNeRV 29.93 33.32 34.86 35.61 36.06 36.35
Boost 34.90 37.35 38.12 38.62 38.70 39.03
Ourts 37.80 38.62 39.00 39.18 39.31 39.40
We evaluate the effectiveness of the proposed SPP-based video representation method across four benchmark datasets: Bunny [32], DAVIS [33], MCL-JCV [34], and UVG [35]. Our method is compared with existing implicit neural video representation baselines, including NeRV, HNeRV, and Boosting-HNeRV (“Boost”). The Bunny dataset consists of a single sequence with a resolution of $720\times 1280$ and 132 frames. To ensure stable training, the frames are cropped to $640\times 1280$. This adjustment is made because HNeRV failed to converge when trained on high-resolution input, as observed in the “Bosphorus” sequence in Table III. The DAVIS dataset consists of 50 natural video sequences with resolution $1080\times 1920$ (cropped $640\times 1280$). Each sequence is relatively short (25 to 104 frames). Similarly, MCL-JCV contains 30 sequences at $1080\times 1920$ (cropped to $640\times 1280$), with lengths ranging from 120 to 150 frames. The UVG dataset comprises 7 long sequences at full HD resolution ($1080\times 1920$), each containing either 300 or 600 frames.
We adjust the stride configuration in the decoder to match the spatial resolution: the stride list is set to [5, 4, 2, 2, 2] for $640\times 1280$, and [5, 3, 2, 2] for $1080\times 1920$. The number of patches is fixed at $N=4$, and we empirically set the loss weighting coefficients to $\alpha=42$ and $\beta=18$. The default model size is 1.5M parameters, except for the UVG dataset at $1080\times 1920$ resolution, where a 3.0M parameter model is used. For quantitative evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Multi-Scale Structural Similarity (MS-SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). In the video compression setting, we use the UVG dataset with a resolution of $1080\times 1920$ and subsample each sequence to 60 or 120 frames. Each video is encoded into multiple models of varying sizes by adjusting the network width. We apply quantization and entropy coding to these models to simulate real-world compression, enabling rate-distortion (RD) analysis.
Quantitative results on DAVIS and MCL-JCV are summarized in Tables I and II, respectively. These tables show the average reconstruction quality across each datasets. Table III reports PSNR and MS-SSIM scores for each sequence in the UVG dataset. Our method consistently achieves higher scores in most sequences, demonstrating superior reconstruction quality compared to baseline methods. In particular, Table IV presents PSNR values at each training epoch for the Bunny dataset, demonstrating not only improved reconstruction performance, but also faster convergence compared to competing methods. Qualitative comparisons are shown in Fig. 4, which visualizes representative frames from DAVIS and MCL-JCV. Our method better preserves sharp edges and spatial coherence across a variety of scenes. Finally, Fig. 5 presents the rate-distortion curves obtained from UVG. The proposed method achieves superior compression performance relative to the baselines. Note that HNeRV is excluded from the average curve due to repeated convergence failures during training.
Figure 5: Compression result on UVG dataset.
V Conclusion
We propose a neural video representation method based on structure-preserving patches (SPPs). In our approach, each video frame of a video is rearranged into spatially consistent patch images, enabling reconstruction that preserves global structure. This strategy mitigates the degradation commonly seen in conventional patch-based decoding. To facilitate smooth reconstruction, we designed a global-to-local decoder architecture. Early decoding layers are conditioned on the temporal index to model the global structure, while later layers incorporate patch-specific information to enhance local details. This separation allows the network to balance structural coherence and fine-grained fidelity effectively. Experimental results demonstrate that our method achieves higher reconstruction quality than existing NeRV-based baselines. These findings highlight the potential of structure-aware patch design for improving the fidelity and efficiency of neural video representations. In future work, we plan to explore more flexible adaptive patch layouts to further enhance spatial coherence and compression performance.